# Constructing Benchmarks and Interventions for Combating Hallucinations in LLMs

### v1

@@ Appendix E Hyperparams for Intervention
We aim to determine the optimal hyperparameters through validation and subsequently apply these parameters to the test set. We will select based on the average score on the classification accuracy on the validation set (200 total) using one seed. We optimized for the classification accuracy since this is the criterion according to which the data is labeled.
For the dynamic version, we will test the TD threshold parameter (TD = 0.65 / 0.7), while for the static version, we will explore the number of layers and heads to intervene on (5 layers and 50 heads / 3 heads and 30 heads).
The best hyperparameters for Llama2 on both DisentQA and TriviaQA are TD = 0.65 for the dynamic intervention and 5 layers and 50 heads for the static version. Full results are provided in Tables 5, 6, 7 and 8.
We also found that for the Goat model, the best TD is 0.65 for the dynamic intervention on both datasets. The full results are presented in Tables 9 and 10.
| component | score (TD=0.65) | score (TD=0.7) |
| Attention | 60.5 | 59.0 |
| MLP | 56.5 | 54.0 |
| Heads | 60.0 | 51.0 |
| Residual | 62.5 | 61.0 |
| None | 50.0 | 50.0 |
| Average | 57.9 | 55.0 |
| component | score (3 layers) | score (5 layers) |
| Attention | 53.0 | 63.0 |
| MLP | 49.5 | 52.0 |
| Heads | 50.0 | 51.5 |
| Residual | 83.0 | 83.5 |
| None | 50.0 | 50.0 |
| Average | 57.1 | 60.0 |
| component | score (TD=0.65) | score (TD=0.7) |
| Attention | 55.5 | 50.0 |
| MLP | 48.5 | 50.0 |
| Heads | 51.0 | 50.0 |
| Residual | 66.5 | 55.5 |
| None | 50.0 | 50.0 |
| Average | 54.3 | 51.1 |
| component | score (3 layers) | score (5 layers) |
| Attention | 58.0 | 56.0 |
| MLP | 46.5 | 48.5 |
| Heads | 49.5 | 50.0 |
| Residual | 68.5 | 82.0 |
| None | 50.0 | 50.0 |
| Average | 54.5 | 57.3 |
| component | score (TD=0.65) | score (TD=0.7) |
| Attention | 63.5 | 62.0 |
| MLP | 62.0 | 61.5 |
| Heads | 65.0 | 61.5 |
| Residual | 62.0 | 64.5 |
| None | 50.0 | 50.0 |
| Average | 60.5 | 59.9 |
| component | score (TD=0.65) | score (TD=0.7) |
| Attention | 54.0 | 49.5 |
| MLP | 57.0 | 50.5 |
| Heads | 48.5 | 49.5 |
| Residual | 70.5 | 50.0 |
| None | 50.0 | 50.0 |
| Average | 56.0 | 49.9 |
@@ Appendix F Additional Results
In this section, we show a few additional results to the ones in Section 6.
@@ F.1 Generation Examples of Static and Dynamic Intervention
To complete the picture from Section 6.1 we show a few more examples of intervention on the residual using dynamic intervention vs. static intervention on Llama2-7B in Table 11 and 12. We can see that the impact of the static intervention has made the text much less comprehensive. The impact on the DisentQA dataset is much stronger than on the TriviaQA dataset, where even the dynamic intervention creates a less comprehensive text.
@@
[…]
F.2 Additional Pre-training vs. Finetuing Results
@@
| Setting | DisentQA | TriviaQA | |||||
| component | direction type | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity |
@@
| Setting | DisentQA | TriviaQA | |||||
| component | direction type | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity |
| Attention | mass-mean | ||||||
@@ F.4 Results on Weighted Dataset
Here are the results of dynamic intervention with on a weighted-labeled dataset. This dataset comprises the test set consisting of 200 hallucinate examples, 200 grounded examples, and 200 other examples, weighted by the count of each label in the full dataset. This approach aids in estimating how intervention will impact a generally distributed dataset. The results are depicted in Figures 16 and 17. These findings align with the results obtained on a dataset evenly split between hallucination and grounded, as presented in the main paper, albeit with weaker trends.
F.5 Generalization of Detection
So far we have performed each experiment within a given dataset and model. Do hallucination detectors generalize across datasets or models? In Figure 18(a), we employ a classifier trained on Goat-7B and evaluate its performance when detecting hallucinations from Llama2-7B states, on the DisentQA dataset (GoatLlama2; DisentQA). Interestingly, the Goat classifier consistently outperforms random chance across all layers, suggesting a notable correlation between the inner states of the two models. These results correspond to related works (Wu et al., 2020). However, the results are much lower than a detector trained and evaluated on Llama2 states (Llama2Llama2; 9(a)), indicating that cross-model generalization is limited.
Similar to these findings, we observe a comparable outcome on the Llama2 TriviaQA setting in Figure 19. Additionally, the results of using the Llama2 model to learn a classifier for the Goat model, which is similar to the results in the other direction, also show accuracy above random, are in Figure 20. Lastly, in Figure 21, we present the results on the Goat model using each dataset to detect hallucinations on the other dataset.

### v2

@@ Appendix E Hyperparams for intervention
We aim to determine the optimal hyperparameters through validation and subsequently apply these parameters to the test set. We select based on the average score on the classification accuracy on the validation set (200 total) using one seed. We optimized for the classification accuracy since this is the criterion according to which the data is labeled.
For the dynamic version, we test the TD threshold parameter (TD = 0.65 / 0.7), while for the static version, we explore the number of layers and heads to intervene on (5 layers and 50 heads / 3 heads and 30 heads).
The best hyperparameters for Llama2 on both DisentQA-WACK and TriviaQA-WACK are TD = 0.65 for the dynamic intervention and 5 layers and 50 heads for the static version.
We also found that for the Goat model, the best TD is 0.65 for the dynamic intervention on both datasets.
@@ Appendix F Additional results
In this section, we show a few additional results to the ones in Section 6. All the results here and in the main paper ran on three seeds (in our code we specifically used 100, 200, None seeds).
@@ F.1 Generation examples of static and dynamic intervention
To complete the picture from Section 6.1 we show a few more examples of intervention on the residual using dynamic intervention vs. static intervention on Llama2-7B in Table 5 and 6. We can see that the impact of the static intervention has made the text much less comprehensive. The impact on the DisentQA-WACK dataset is much stronger than on the TriviaQA-WACK dataset, where even the dynamic intervention creates a less comprehensive text.
@@
F.2 Additional pre-training vs. finetuning results
@@
| Setting | DisentQA-WACK | TriviaQA-WACK | |||||
| Classification | Generation | Classification | Generation | ||||
| Accuracy | Accuracy | Wiki ppl | Accuracy | Accuracy | Wiki ppl | ||
@@
| Setting | DisentQA-WACK | TriviaQA-WACK | |||||
| Classification | Generation | Classification | Generation | ||||
| Accuracy | Accuracy | Wiki ppl | Accuracy | Accuracy | Wiki ppl | ||
| Attention | MM | ||||||
@@ F.4 Results on weighted dataset
Here are the results of the basic configuration (see Section 5) on a weighted-labeled dataset. This dataset comprises the test set consisting of 200 hallucination examples, 200 grounded examples, and 200 other examples (chosen from a random of 1000 else labeled examples), weighted by the count of each label in the full dataset. This approach aids in estimating how the intervention impact a generally distributed dataset. The results are depicted in Figures 20 and 21. These findings align with the results obtained on a dataset evenly split between hallucination and grounded, as presented in the main paper, albeit with weaker trends.